OpenStudy (anonymous):

what is the pattern -4 -1 2

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\text{u can check the common difference here}\\~\\~\\ -1-(-4)=3\\ ~~~~2-(-1)=3\\~\\~\\~~\\ \text{so this is arithmatic sequence with common difference 3}\\~\\~~\\~\\~~\\ -4,-1,2,5,8,11........}\end{align}\)